OpenStudy (anonymous):

-8x-8y=-8 -9x-9y=9 is the answer no solution, (4,-1), (7,-1), or (4,1)?

OpenStudy (anonymous):

Simplify the equation for line 1 to get it into our y = mx + b format: --8x - 8y = -8 -8y = -8x - 8 Divide each side of the equation by -8 to isolate y: -8y -8 = -8x - 8 -8 Simplifying and evaluating, we have: y = -1x + 1 Therefore, the slope of line equation 1 = -1 Simplify the equation for line 2 to get it into our y = mx + b format: --9x - 9y = 9 -9y = -9x + 9 Divide each side of the equation by -9 to isolate y: -9y -9 = -9x + 9 -9 Simplifying and evaluating, we have: y = -1x - 1 Therefore, the slope of line equation 2 = -1 Since the slope of both lines = -1, the lines are parallel and never cross The 2 lines are parallel. NO SOLUTION